Calculate the volume of the folllowing at S,T.P ?
a. 3.2g of O2;
b.51g of H2S;
c.55g of CO2;
1
Expert's answer
2020-07-07T14:35:11-0400

One mole of each gas occupies 22.4 L. As a result, the volume of gases can be calculated using the following equation:

V = n × 22.4 L/mol = (m / Mr) / 22.4 L/mol

where V - volume, n - number of moles, m - mass, Mr - molecular weight.

From here:

a) Mr(O2) = 32 g/mol

V(O2) = (3.2 g / 32 g/mol) × 22.4 L/mol = 2.24 L

b) Mr(H2S) = 34 g/mol

V(H2S) = (51 g / 34 g/mol) × 22.4 L/mol = 33.6 L

c) Mr(CO2) = 44 g/mol

V(CO2) = (55 g / 44 g/mol) × 22.4 L/mol = 28 L


Answer: a) 2.24 L, b) 33.6 L, c) 28 L